OpenStudy (anonymous):

How do I factor 15p^2-26p+11?

myininaya (myininaya):

lets find two factors of 15(11) that add up to be -26 (-15)(-11) -15+(-11)=-26 so 15p^2-26p+11 = 15p^2-15p-11p+11 =15p(p-1)-11(p-1) =(p-1)(15p-11)

OpenStudy (anonymous):

Thank you